OpenStudy (anonymous):

Which expression is equivalent to tan (-x)cos^2x? A)-sin xcos x B)sin xcos x C) tan x + sin^2 D) tan x - sin^2 E) -tan x-sin^2x

OpenStudy (mathstudent55):

Use the identity \(\sin^2 x + \cos^2 x = 1\). Solve it for \(\cos^2 x\). Then substitute it into the given expression. What do you get?